Grade 12th pass10 grade science

78. An electric iron draws a cu,.ent of 0.6 A when the voltage is 100 volt Calculate the amount of electric charge flowing through it in one hour.

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Answer: Explanation: As we know, Q = I × t ∴ Q = 0.6 × 3600 = 2160 C